OpenStudy (anonymous):

Use substitution to determine which of the following points is a solution to the standard form equation below 5x-2y=10

OpenStudy (studygurl14):

I think some info is missing...where are the points?

OpenStudy (anonymous):

these are the points: -1,5. 1,5. 0,-5. 0,5

OpenStudy (studygurl14):

Okay, so all you have to do is plug in each answer choice into the equation until you get one that works. Tha'ts the answwer

OpenStudy (studygurl14):

The first answer choice: (-1, 5) 5x - 2y = 10 5(-1) -2(5) = 10 -5 - 10 = 10 This does not work, so A can't be the answer. See what I did?

OpenStudy (anonymous):

yes i do but what one is the answer?

OpenStudy (studygurl14):

You should do it yourself. I showed you how, please do the rest. I don't want to just give you the answer

OpenStudy (anonymous):

but i cant do it that fast and my test is timed:(

OpenStudy (studygurl14):

This is a test? Now I'm not going to give you the answer even more so. Sorry.

OpenStudy (anonymous):

i only have 15 more mins to answer 10 questions

OpenStudy (anonymous):

help pleaseeee!

OpenStudy (studygurl14):

Fine. I'll help because I'm so nice... The answer is C

OpenStudy (anonymous):

Use substitution to determine which of the following points is a solution to the standard form equation below. 2y=5 help again please

OpenStudy (studygurl14):

what are the points?

OpenStudy (anonymous):

10,(5/2), 5(2/5), (5/2)(-5/2), (5/2)(-2/5).

OpenStudy (studygurl14):

A

OpenStudy (anonymous):

Is the equation below written in standard form? If not, select which explanation best applies to why the equation is not written in standard form. -2x+3y=-5

OpenStudy (anonymous):

the requirement that A >_ 0 is not satisfied the requirement that A and B are not both zero is not satisfied A, B, and C do not have a greatest common factor of 1 the equation is written in standard form

OpenStudy (studygurl14):

No, I don't think it is in standard form. because stand form is y = mx + b, I'm not sure on this one tho

OpenStudy (studygurl14):

I don't know about the other one

OpenStudy (anonymous):

why isnt it writtin in stand out of those chocies?